OpenStudy (anonymous):

Q: K > L > M > N are positive integers such that KM + LN = (K + L - M + N)(-K + L + M + N). Prove that KL + MN is not prime.

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

since KL+MN=(K+M-L+N)(-K+M+L+N), then we know that will not only be divisible by one and itself. It will also be divisible by K+M-L+N and -K+M+L+N since it is also a product of these too.

OpenStudy (anonymous):

How did you get KL+MN=(K+M-L+N)(-K+M+L+N)?

OpenStudy (anonymous):

No \[(K+M-L+N) (-K+L+M+N)=\\-K^2+2 K L-L^2+M^2+2 M N+N^2 \]

OpenStudy (anonymous):

@eliassaab what do you mean?

OpenStudy (anonymous):

I mean \[ KL+MN\ne(K+M-L+N)(-K+M+L+N) \]

OpenStudy (anonymous):

because KM+LN=(K+L-M+N)(-K+M+N+L) so I've mistaken that for KL+MN=(K+M-L+N)(-K+M+L+N)

OpenStudy (anonymous):

sorry :))

OpenStudy (anonymous):

since we know KM + LN is not prime, if we can show it shares a common factor with KL + MN i think that will do it

OpenStudy (anonymous):

@eigenschmeigen how do I do it?

OpenStudy (anonymous):

im not sure, i havent really attempted anything

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

@nikvist

OpenStudy (anonymous):

i stuck the first equation in wolfram, says there aren't any positive integer solutions what does this mean? are we still able to prove this? http://www.wolframalpha.com/input/?i=ab+%2B+cd+%3D+%28a+%2B+c+-+b+%2B+d%29%28-a+%2B+c+%2B+b++%2B+d%29

OpenStudy (anonymous):

@aron_west

OpenStudy (anonymous):

Yes, finally someone is replying!

OpenStudy (anonymous):

@eliassaab , in expanding the right side of "KM + LN = (K + L - M + N)(-K + L + M + N)", i got 2KM + 2LN terms along with some other terms.... anybody got that too?

OpenStudy (anonymous):

Me, Me ...MEEEHH! I got that as well. \[KM + LN = K^2 + M^2 - L^2 - N^2\]

OpenStudy (anonymous):

yah, thats what i got...

OpenStudy (anonymous):

I posted this problem on m.se, none got the solution there. I think something is not right.

OpenStudy (anonymous):

@eigenschmeigen That means the problem has something not right with it.

OpenStudy (anonymous):

Nothing is wrong with the problem, it's an IMO problem or Non-solvable by commons like me. http://imo.wolfram.com/problemset/IMO2001_solution6.html